# Blown up by an equilateral: Poncelet triangles about the incircle and their degeneracies

### v1

@@ Abstract.
We tour several harmonious Euclidean properties of Poncelet triangles inscribed in an ellipse and circumscribing the incircle. We also show that a number of degenerate behaviors are triggered by the presence of an equilateral triangle in the family.
@@ 1. Introduction
Poncelet’s porism is a 1d family of $n$-gons with vertices on an ‘outer’ conic $\mathcal{E}$ and with sides tangent to an ‘inner’ conic $\mathcal{E}_{c}$ (also called the ‘caustic’). For such a porism to be admitted, $\mathcal{E}$ and $\mathcal{E}_{c}$ must be positioned in $\mathbb{R}^{2}$ so as to satisfy ‘Cayley’s condition’ [7]. While the porism is a projective statement – most conic pairs can be projected to two circles where the ‘Poncelet map’ is linearized, – we have found that Poncelet triangles are a wellspring of interesting Euclidean phenomena.
Consider when either conic can be a circle or a generic ellipse. Referring to Figure 1, we obtain the following four combinations:
  1. (1)
    Both conics are circles: this is Chapple’s porism [3]. Loci loci of its triangle centers, e.g., incenter $X_{1}$, barycenter $X_{2}$, circumcenter $X_{3}$, orthocenter $X_{4}$ (the $X_{k}$ notation, used here, follows [18]), have been studied in [13, 14, 21, 26]. See Figure 2 (left).
  2. (2)
    Only $\mathcal{E}$ is a circle. This system, always an affine image of (4), lends itself to a convenient symmetric parametrization, explained below.
  3. (3)
    Only $\mathcal{E}_{c}$ is a circle: the present study is a tour of its manifold harmonious properties.
  4. (4)
    No restriction:though Euclidean phenomena are sparser, some are still discernible, e.g., the locus of certain centers are always conics, the locus of barycenter $X_{2}$ and orthocenter $X_{4}$ always homothetic to $\mathcal{E}$ (see below), etc.
Another special case is that a triangle family interscribed between two confocal ellipses (the ‘elliptic billiard’), shown in Figure 2 (right). The loci of all four ‘Greek centers’ of the triangle – incenter, barycenter, circumcenter, orthocenter – sweep ellipses, as proved in [8, 11, 29].
@@
For case (3), studied here, we start by identifying triangle centers whose loci are conics. In [17] we show that the locus of a triangle center which is a fixed linear combination of the barycenter and the orthocenter is a sufficient condition for that center to sweep a conic. Nevertheless, other centers may still sweep conics for some choices of conic pairs. Still lacking a predictive theory for those, we use numerical methods to pinpoint these, as in [11].
We find many curious dependencies between the shape and foci of certain loci and the specifications of both Poncelet conics. A surprising find has been that the presence of a single equilateral triangle in the family triggers degeneracies in certain loci.
@@ 1.1. Article structure
  • Section 3: we characterize remarkable properties of loci of certain triangle centers, including the barycenter, circumcenter, orthocenter, Euler center, and the Gergonne and Nagel points, labeled in [18] as $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{7}$, and $X_{8}$, respectively. The behaviors include:
@@
    • The ‘pinning’ or ‘railing’ of the foci of certain centers to the axes of $\mathcal{E}$ or to the caustic center $C$.
@@
  • Section 4: A series of curious, ‘degenerate’ locus phenomena occurs if the Poncelet family contains an equilateral triangle. This occurs when $C$ is on a special ellipse $\mathcal{E}_{eq}$ which is the locus of the centroids of all equilaterals inscribed in $\mathcal{E}$, and derived in [34]. We prove the following phenomena:
@@
      The Feuerbach point $X_{11}$ becomes stationary on the incircle, and the reflection of the incenter about it ($X_{80}$) is pinned to a point on $\mathcal{E}$.
@@
      The aspect ratio of the loci of the circumcenter $X_{3}$ and the Gergonne point $X_{7}$ are invariant over all $C$ on $E_{eq}$.
@@
      The locus of $X_{59}$, the ‘isogonal conjugate’ of $X_{11}$ (a type of inversive transformation [36]), normally a high-degree curve, ‘tames down’ to an ellipse which touches $\mathcal{E}$ at a special point.
  • Section 5: the harmonious envelopes of both the circumcircle and of the (incircle,circumcircle) radical axis are studied. The former is shown to have two circular boundaries centered on the foci of the $X_{3}$ locus, and the latter is shown to be always a conic, which degenerates to a parabola if $C$ is on the boundary $\mathcal{E}_{eq}$. We conjecture that the envelope of said radical axis is a conic if and only if the Poncelet family circumscribes a circle.
  • Section 6: by comparing several Poncelet triangle families, we conjecture that the loci of $X_{7}$ and $X_{8}$ can only be ellipses if the caustic is a circle or the pair is confocal.
  • Section 7: we describe four special triangle porisms about a fixed incircle, that keep key triangle centers stationary: $X_{1}$, $X_{2}$, $X_{4}$, and $X_{7}$. We also identify their conservations, some of which generalize to $n>3$ Poncelet porisms.
@@
  • Appendix B: we categorize several dozen triangle centers according to one of six remarkable ‘behaviors’: E-homothety, axis-alignment, focus on $C$, major (resp. minor) axis through $C$, and circularity.
  • Appendix C: we analyze the circle-inscribed ‘contact’ family, showing that some of its phenomena are closely related to those of the original family.
  • Appendix D: Long expressions for the semiaxes of the locus of the Gergonne point $X_{7}$.
  • Appendix E compiles most symbols used in the article.
In [6, p.210] it is said that an experimental discovery approach “offers the possibility in some instances of formalizing the inductive leaps that the mathematical mind takes when confronted with what seems to be, logically speaking, incomplete evidence”. This has certainly been the case for us as most phenomena have been discovered via combination. We include include links to animations of the relevant phenomena in some figure captions as well as links to experiments with our interactive tool [5].
@@ 2.2. Symmetric parametrization
In order to generalize this parameterization for any Poncelet family of triangles interscribed between two ellipses, we simply apply an affine transformation that takes the unit circle centered at the origin to the desired outer ellipse. Without loss of generalization, we can assume that the outer ellipse is also centered at the origin and its major axis aligned with the real axis of $\mathbb{C}$.
As before, let $\mathcal{E}$ and $\mathcal{E}_{c}$ be the outer and inner ellipses of a Poncelet family of triangles and let $a$ and $b$ be the lengths of the semi-major and semi-minor axes of $\mathcal{E}$, respectively, where $\mathcal{E}$ is axis-aligned and centered at the origin. We can then use the transformation $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ defined by $T(x,y)=(ax,by)$ to transform the unit circle into $\mathcal{E}$. With a slight abuse of notation, this transformation can also be represented in the complex plane, $T:\mathbb{C}\rightarrow\mathbb{C}$, as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\overline{z}$
@@
Let $P:\mathbb{C}^{3}\rightarrow\mathbb{C}$ be a symmetric rational function. Our objective is to parameterize the point $P(z_{1},z_{2},z_{3})$. Since $P$ is symmetric on its $3$ inputs, the function $P(T(\cdot),T(\cdot),T(\cdot)):\mathbb{C}^{3}\rightarrow\mathbb{C}$ must also be symmetric on its $3$ inputs. Restricting $T$ to the unit circle $\mathbb{T}$, we can write $T:\mathbb{T}\rightarrow\mathcal{E}$ as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\frac{1}{z}$, which is itself a rational function. Thus, $P(z_{1},z_{2},z_{3})=P(T(z_{1}^{\prime}),T(z_{2}^{\prime}),T(z_{3}^{\prime}))$ is a symmetric rational function on $z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}$. Write $P(z_{1},z_{2},z_{3})=\frac{P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}{P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}$ where $P_{1}$ and $P_{2}$ are symmetric polynomials. By the Fundamental Theorem of Symmetric Polynomials, we can express $P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ and $P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ as polynomials in $\sigma_{1},\sigma_{2},\sigma_{3}$ where $\sigma_{1}:=z_{1}^{\prime}+z_{2}^{\prime}+z_{3}^{\prime}$, $\sigma_{2}:=z_{1}^{\prime}z_{2}^{\prime}+z_{2}^{\prime}z_{3}^{\prime}+z_{3}^{\prime}z_{1}^{\prime}$, and $\sigma_{3}:=z_{1}^{\prime}z_{2}^{\prime}z_{3}^{\prime}$. Substituting the symmetric parameterization from Theorem Theorem 1, we finally arrive at a parameterization for $P(z_{1},z_{2},z_{3})$ with $\lambda\in\mathbb{T}$ as the parameter, as desired.
@@ 3.1. Loci of X(k), k=2,4,7
to do ∎
@@ 3.2. Loci of X(k), k=3,5,8
to do ∎
@@ 4.5. Line X(36)
Figure 16. Over all $C=X_{1}$ on $\mathcal{E}_{eq}$ (dashed green), the envelope of the line locus of $X_{36}$ is an oval of degree 6 (dashed orange).
@@ 4.6. Elliptic X(59)
Over $\mathcal{T}_{g}$, the locus of $g(P)$ is a conic. If $P$ is on $\mathcal{E}_{c}$ the locus is an ellipse interior to $\mathcal{E}$ and touching it at a point.
Figure 18. Results from [10]: left: over generic Poncelet triangles, the locus of the isogonal conjugate $g(P)$ of a fixed point $P$ is a conic (purple); right: if $P$ is on the caustic, the locus of $g(P)$ is an ellipse that touches $\mathcal{E}$ at one point.
@@ 5.2. Radical axis
Developing the calculations it follows that the envelope is parametrized by:
@@
Here all constants are long expressions involving the variables $(a,b,c,x_{c},y_{c}).$ Eliminating the variables we obtain that the envelope is given implicitly by a quadratic equation and so it is a conic.
to do. ∎
@@
The orthic axis, called $\mathcal{L}_{3}$ in [36] (resp. anti-orthic axis, $\mathcal{L}_{1}$) is the radical axis of the circumcircle and the nine-point (resp. Bevan) circle [36, Radical line], the latter is centered on $X_{5}$ (resp. $X_{40}$) and has radius $R/2$ (resp. $2R$). Experimentation shows:

### v2

@@ Abstract.
We tour several Euclidean properties of Poncelet triangles inscribed in an ellipse and circumscribing the incircle. We also show that a number of degenerate behaviors are triggered by the presence of an equilateral triangle in the family.
@@ 1. Introduction
Poncelet’s porism is a 1d family of $n$-gons with vertices on a first conic $\mathcal{E}$ and with sides tangent to a second conic $\mathcal{E}_{c}$ (also called the ‘caustic’). For such a porism to exist, $\mathcal{E}$ and $\mathcal{E}_{c}$ must be positioned in $\mathbb{R}^{2}$ so as to satisfy ‘Cayley’s condition’ [7]. While the porism is projectively invariant – a conic pair is the projective image of two circles where the ‘Poncelet map’ is linearized, – we have found that Poncelet porisms of triangles ($n=3$) are a wellspring of interesting phenomena involving the dynamical geometry of classical objects associated with the triangle (centers, circles, lines/axes, etc.).
An emblematic case, shown in Figure 2 (right) is when when $\mathcal{E},\mathcal{E}_{c}$ are confocal ellipses. In such a porism, – also known as the elliptic billiard – the perimeter is conserved [35], and the loci of all four ‘Greek centers’ of the triangle – incenter, barycenter, circumcenter, orthocenter – sweep ellipses [8, 11, 29]. Significantly, neither foci, perimeter, nor triangle centers are projectively equivariant.
@@
Guided by experiment, we have found that certain types of $\mathcal{E},\mathcal{E}_{c}$ pairs form equivalent classes of Euclidean phenomena. One criterion is whether the pair is concentric and axis-aligned or not. Relevant for the present study, is a subdivision of triangle porisms based on whether either Poncelet conic is a circle or not. Referring to Figure 1, we obtain the following four combinations:
  1. (1)
    Both conics are circles: this is Chapple’s porism [3]. Loci of its triangle centers, e.g., incenter $X_{1}$, barycenter $X_{2}$, circumcenter $X_{3}$, orthocenter $X_{4}$ (the $X_{k}$ notation, used here, follows [18]), have been studied in [13, 14, 21, 26]. See Figure 2 (left).
  2. (2)
    Only $\mathcal{E}$ is a circle. This system, always an affine image of (4), lends itself to a symmetric parametrization which is at the basis of our analysis. Aspects of its geometry and phenomena are analyzed in Appendix C.
  3. (3)
    Only $\mathcal{E}_{c}$ is a circle: here we will tour its manifold Euclidean properties.
  4. (4)
    Neither is circular: Euclidean phenomena are sparser, some are still discernible, e.g., the locus of certain centers are always conics, the locus of barycenter $X_{2}$ and orthocenter $X_{4}$ always homothetic to $\mathcal{E}$ (see below), etc.
For case (3), we start by identifying triangle centers whose loci are conics. In [17] we show that the locus of a triangle center which is a fixed linear combination of the barycenter and the orthocenter is a sufficient condition for that center to sweep a conic. Nevertheless, other centers may still sweep conics for some choices of conic pairs. Still lacking a predictive theory for the latter, we use numerical methods to pinpoint their conic loci, as in [11].
Below we describe many curious dependencies between the shape and foci of certain loci and the geometry of either Poncelet conic. A surprising find has been that the presence of a single equilateral triangle in the family triggers a host of harmonious degeneracies.
@@ 1.1. Article structure
  • Section 3: we characterize remarkable properties of loci of certain triangle centers, including the barycenter, circumcenter, orthocenter, Euler center, and the Gergonne and Nagel points, denoted $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{7}$, and $X_{8}$, respectively, after [18]. The behaviors include:
@@
    • The ‘railing’ of the foci of certain centers to the axes of $\mathcal{E}$ or to the caustic center $C$.
@@
  • Section 4: A series of curious, ‘degenerate’ phenomena occurs if the Poncelet family contains an equilateral triangle. This is tantamount to $C$ lying on a special ellipse $\mathcal{E}_{eq}$ which is the locus of the centroids of all equilaterals inscribed in $\mathcal{E}$, and derived in [34]. The following facts arise:
@@
      The Feuerbach point $X_{11}$ becomes stationary on the incircle, and the reflection of the incenter about it ($X_{80}$) is a fixed point on $\mathcal{E}$.
@@
      The aspect ratio of the loci of the circumcenter $X_{3}$ and the Gergonne point $X_{7}$ are invariant over all $C$ on $\mathcal{E}_{eq}$.
@@
      The locus of $X_{59}$, the ‘isogonal conjugate’ of $X_{11}$ (a type of inversive transformation [37]), normally a high-degree curve, collapses to an ellipse touching $\mathcal{E}$ at a special point.
  • Section 5: the envelopes of both the circumcircle and of the (incircle,circumcircle) radical axis are studied. The former is shown to be the union of two circles centered on the foci of the $X_{3}$ locus; the latter is shown to be always a conic, which degenerates to a parabola if $C$ is on the boundary $\mathcal{E}_{eq}$. We conjecture that either envelope has a conic component if and only if the Poncelet family circumscribes a circle.
  • Section 6: by comparing several Poncelet triangle families, we conjecture that the loci of the Gergonne point $X_{7}$ and the Nagel point $X_{8}$ are conics if and only if (i) the caustic is a circle or (ii) the pair is confocal.
  • Section 7: we describe four special triangle porisms when $\mathcal{E}_{c}$ is a circle, that keep key triangle centers stationary: $X_{1}$ (on a focus of $\mathcal{E}$), $X_{2}$, $X_{4}$, and $X_{7}$. We also identify their conservations, some of which generalize to $n>3$ Poncelet porisms.
@@
  • Appendix B: the loci of several dozen triangle centers are categorized according to one of six remarkable ‘behaviors’: $\mathcal{E}$-homothety, axis-alignment, focus on $C$, major (resp. minor) axis through $C$, and circularity.
  • Appendix C: a geometric analysis of circle-inscribed Poncelet triangles, or case (2) above; we show that phenomena here are closely related to those in (3).
  • Appendix D: computational expressions for the semiaxes of the locus of the Gergonne point $X_{7}$.
  • Appendix E: a table with most symbols used in the article.
An experimental discovery approach “offers the possibility in some instances of formalizing the inductive leaps that the mathematical mind takes when confronted with what seems to be, logically speaking, incomplete evidence” [6, p.210]. This has certainly been the case for us as most phenomena have been discovered via simulation. In some figure captions we include links to videos as well as interactive animations with a custom-built tool [5].
@@ 2.2. Symmetric parametrization
In order to generalize this parameterization for any Poncelet family of triangles interscribed between two ellipses, we simply apply an affine transformation that takes the unit circle centered at the origin to the desired outer ellipse. Without loss of generality, we can assume that the outer ellipse is also centered at the origin and its major axis is aligned with the real axis of $\mathbb{C}$.
As before, let $\mathcal{E}$ and $\mathcal{E}_{c}$ be the outer and inner ellipses of the Poncelet family of triangles and let $a$ and $b$ be the lengths of the semi-major and semi-minor axes of $\mathcal{E}$, respectively, where $\mathcal{E}$ is axis-aligned and centered at the origin. We can then use the transformation $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ defined by $T(x,y)=(ax,by)$ to transform the unit circle into $\mathcal{E}$. With a slight abuse of notation, this transformation can also be represented in the complex plane, $T:\mathbb{C}\rightarrow\mathbb{C}$, as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\overline{z}$. Moreover, we have $T^{-1}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ acting as $T^{-1}(x,y)=(x/a,y/b)$ or alternatively, $T^{-1}:\mathbb{C}\rightarrow\mathbb{C}$ acting on $\mathbb{C}$ as $T^{-1}(z)=\frac{(1/a+1/b)}{2}z+\frac{(1/a-1/b)}{2}\overline{z}$.
@@
Let $P:\mathbb{C}^{3}\rightarrow\mathbb{C}$ be a symmetric rational function. Our objective is to parameterize the point $P(z_{1},z_{2},z_{3})$. Since $P$ is symmetric on its $3$ inputs, the function $P(T(\cdot),T(\cdot),T(\cdot)):\mathbb{C}^{3}\rightarrow\mathbb{C}$ must also be symmetric on its $3$ inputs. Restricting $T$ to the unit circle $\mathbb{T}$, we can write $T:\mathbb{T}\rightarrow\mathcal{E}$ as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\frac{1}{z}$, which is itself a rational function. Thus, $P(z_{1},z_{2},z_{3})=P(T(z_{1}^{\prime}),T(z_{2}^{\prime}),T(z_{3}^{\prime}))$ is a symmetric rational function on $z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}$. Write $P(z_{1},z_{2},z_{3})=\frac{P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}{P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}$ where $P_{1}$ and $P_{2}$ are symmetric polynomials. By the Fundamental Theorem of Symmetric Polynomials, we can express $P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ and $P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ as polynomials in $\sigma_{1},\sigma_{2},\sigma_{3}$ where $\sigma_{1}:=z_{1}^{\prime}+z_{2}^{\prime}+z_{3}^{\prime}$, $\sigma_{2}:=z_{1}^{\prime}z_{2}^{\prime}+z_{2}^{\prime}z_{3}^{\prime}+z_{3}^{\prime}z_{1}^{\prime}$, and $\sigma_{3}:=z_{1}^{\prime}z_{2}^{\prime}z_{3}^{\prime}$. Substituting the symmetric parameterization from Theorem 1, we arrive at a parameterization for $P(z_{1},z_{2},z_{3})$ with $\lambda\in\mathbb{T}$ as the parameter, as desired.
With this method, we have proved many results in [16] and [17] such as:
Theorem 2.
In any Poncelet family of triangles, the loci of the barycenter, circumcenter, orthocenter, nine-point center, and the de Longchamps point are always ellipses. More generally, any fixed linear combination of the barycenter $X_{2}$ and the circumcenter $X_{3}$ traces out an ellipse as its locus.
This method is great for proving results such as this, but not so great for finding explicit formulas for the loci. This is due to the fact that everything depends on the foci $f,g$ of the ellipse $\mathcal{E}_{pre}$ which is generally neither of the ellipses $\mathcal{E},\mathcal{E}_{c}$ that define our Poncelet family. However, particularizing the case where the inner ellipse is a circle will allow us to compute such formulas explicitly from the original data of $\mathcal{E}$ and the circle $\mathcal{E}_{c}$.
From now on, assume that $\mathcal{E}_{c}$ is a circle with center $C=x_{c}+iy_{c}$ where $x_{c},y_{c}\in\mathbb{R}$ and radius $r$ given as in Proposition 2.
Since $T^{-1}$ stretches the real and complex axis by factors of $1/a$ and $1/b$, respectively, $\mathcal{E}_{pre}:=T^{-1}(\mathcal{E}_{c})$ is an axis-aligned ellipse with semi-major axis and semi-minor axis equal to $r/b$ and $r/a$, respectively. In fact, it is a $90^{\circ}$-rotated copy of the outer ellipse $\mathcal{E}$ shrunken by a factor of $r/(ab)$. Therefore its semi-focal length is given by $r\frac{c}{ab}$. Moreover, since ellipse centers are preserved by linear transformations, the center of $\mathcal{E}_{pre}$ is $T^{-1}(C)=x_{c}/a+iy_{c}/b$. Since the major axis of $\mathcal{E}_{pre}$ is parallel to the imaginary axis of $\mathbb{C}$, its foci $f$ and $g$ will be given by $x_{c}/a+iy_{c}/b\pm r\frac{c}{ab}$. In order to use the parameterization from Theorem 1 we explicitly calculate their sum and product as
$f+g=$ $\frac{2x_{c}}{a}+\frac{2y_{c}}{b}$
$fg=$ $\frac{a^{2}+b^{2}}{c^{2}}+\frac{2i}{ab}\left(x_{c}y_{c}+\frac{1}{c}\sqrt{(a^{4}-c^{2}x_{c}^{2})(b^{4}-c^{2}y_{c}^{2})}\right)$
We can now calculate explicit formulas for any Poncelet family of triangles where the inner ellipse is a circle. In particular, we can calculate the loci of the barycenter $X_{2}$ and the circumcenter $X_{3}$, and then use the table in Appendix A to explicitly calculate the loci of many triangle centers that can be conveniently written as a linear combination of $X_{1}$, $X_{2}$, and $X_{3}$.
@@ 3.1. Loci of X(k), k=2,4,7
Symmetric parametrization yields an axis-aligned conic. ∎
@@ 3.2. Loci of X(k), k=3,5,8
Symmetric parametrization yields a concentric, axis-aligned conic. ∎
@@ 4.5. Line X(36)
Figure 16. Over all $C=X_{1}$ on $\mathcal{E}_{eq}$ (dashed green), the envelope of the line locus of $X_{36}$ is a non-conic of degree 6 (dashed orange).
@@ 4.6. Elliptic X(59)
Over $\mathcal{T}_{g}$, the locus of $g(P)$ is a conic. If $P$ is on $\mathcal{E}_{c}$ the locus is an ellipse interior to $\mathcal{E}$ and touching it at a point. If $P$ is on $\mathcal{E}$ the locus is the union of a straight line and the line at infinity (degenerate hyperbola).
An alternative proof to Proposition 17 is to based on the fact that $X_{36}$ and $X_{80}$ are isogonal conjugates [18]:
Corollary 6.
Over $\mathcal{T}^{*}$, the locus of $X_{36}$ is a line since $X_{80}$ is stationary on $\mathcal{E}$, Proposition 15.
Figure 18. Results from [10]: left: over generic Poncelet triangles, the locus of the isogonal conjugate $g(P)$ of a fixed point $P$ is a conic (purple); right: if $P$ is on the caustic, the locus of $g(P)$ is an ellipse that touches $\mathcal{E}$ at one point. Video
@@ 5.2. Radical axis
It follows from the calculations that the envelope is parametrized by:
@@
Here all constants are long expressions involving the variables $(a,b,c,x_{c},y_{c}).$ Eliminating the variables we obtain that the envelope is given implicitly by a quadratic equation and so it is a conic. Algebraic manipulation yields its major axis to be as the one in Proposition 6. ∎
@@
The orthic axis, called $\mathcal{L}_{3}$ in [37] (resp. anti-orthic axis, $\mathcal{L}_{1}$) is the radical axis of the circumcircle and the nine-point (resp. Bevan) circle [37, Radical line], the latter is centered on $X_{5}$ (resp. $X_{40}$) and has radius $R/2$ (resp. $2R$). Experimentation shows: